Solve x^(2)-6x+4=0 by completing the square
Please show work

Answer by checkley79(3341)

x^(2)-6x+4=0
X^2-6X=-4
X^2-6X+9=-4+9
X^2-6X+9=5
(X-3)^2=5 ANS.